### v2

# Subspace Selected Variational Quantum Configuration Interaction with a Partial Walsh Series

Abstract
Estimating the ground-state energy of a quantum system is one of the most promising applications for quantum algorithms. Here we propose a variational quantum eigensolver (VQE) Ansatz for finding ground state configuration interaction (CI) wavefunctions. We map CI for fermions to a quantum circuit using a subspace superposition, then apply diagonal Walsh operators to encode the wavefunction. The algorithm can be used to solve both full CI and selected CI wavefunctions, resulting in exact and near-exact solutions for electronic ground states. Both the subspace selection and wavefunction Ansatz can be applied to any Hamiltonian that can be written in a qubit basis. The algorithm bypasses costly classical matrix diagonalizations, which is advantageous for large-scale applications. We demonstrate results for several molecules using quantum simulators and hardware.
I Introduction
The energy eigenvalue estimation of many-body systems is a computational challenge inspiring continued improvements in both quantum and classical algorithms [1, 2, 40, 9, 36, 39, 48, 51, 41]. In strongly correlated regimes classical calculations can be limited by the growth of the parameter space or many-body entanglement [66, 38, 23, 50]. Quantum algorithms offer asymptotic advantages by natively representing quantum states, but may require large-scale error-correcting or fault-tolerant quantum computers to be competative platforms for routine eigenvalue estimation [33].
While many algorithms have been developed for quantum eigenvalue estimation, [34, 21, 75, 44, 62] we emphasize here variational quantum eigensolvers (VQEs) and quantum selected configuration interaction (QSCI) [1, 9, 49, 31, 1, 41, 77, 76, 27, 74, 67, 6, 54, 32, 12, 57, 63, 18, 40] for electronic Hamiltonians. Currently, noisy intermediate-scale quantum (NISQ) devices make VQE an attractive approach because VQEs require relatively modest quantum resources. Significiant progress has been made developing resource-efficient and physically-motivated VQE Ansätze for molecular Hamiltonians [31, 41, 77, 76, 27, 74, 67, 6, 54]. One drawback of standard VQE Ansätze is the possibility of overparameterization and barren plateaus, which are flat areas of the potential surface that stymie optimization progress [43, 37, 52]. Recently, sample-based or QSCI has shown impressive results for ground-state energy estimation [32, 12, 57, 75, 63, 18]. In QSCI, a wavefunction is prepared on a quantum circuit, the important Slater determinants (SDs) are sampled, and the Hamiltonian is diagonalized classically, where the Hamiltonian is defined within the subspace spanned by the sampled SDs [45, 57, 75, 63, 18, 12]. However, even in the SCI subspace, classical diagonalization of the Hamiltonian can be cost prohibitive, or measurement error can cause the diagonalization of an erroneous Hamiltonian [55].
In this Article, we propose a diagonal Ansatz for VQE, which controls overparametrization of the state and avoids barren plateaus. Notably, the number of one- and two-qubit gates scales linearly with the number of SDs. There are two main components of the algorithm: superposition preparation and Ansatz application. First, on a quantum circuit we prepare a superposition over selected physically-relevant bitstrings. Second, we apply a diagonal Ansatz which effectively encodes the wavefunction coefficients in the appropriate bitstrings. The diagonal operator requires one ancilla qubit, and results in a probabalistic algorithm; however, the success probability is always 50%. Following standard VQE, we estimate the energy of the quantum circuit, and update the coefficients classically. We compute the ground-state energy for a range of molecules with IBMQ’s Torino processor and simulator. Our results provide a path forward for systematically improvable, and non-overparameterized Ansätze for many-body Hamiltonians, with relatively simple and physically motivated quantum circuits.
II Theory and Methods
A molecular electronic Hamiltonian under the Born-Oppenheimer approximation can be expressed as [65],
$\hat{H}=\sum_{mn}K^{m}_{n}\hat{a}^{\dagger}_{m}\hat{a}_{n}+\sum_{mnpq}V^{mn}_{pq}\hat{a}^{\dagger}_{m}\hat{a}^{\dagger}_{n}\hat{a}_{q}\hat{a}_{p},$ (1)
where $\hat{a}$ and $\hat{a}^{\dagger}$ are fermionic annihilation and creation operators, and $K^{m}_{n}$ and $V^{mn}_{pq}$ are one- and two-body integrals respectively. For this Hamiltonian, a general $N$-electron wavefunction, $|\Psi\rangle$, can be written as a linear combination of SDs,
$|\Psi\rangle=\sum_{k=1}^{D}c_{k}|\psi_{k}\rangle,$ (2)
where $D$ is the number of SDs in the expansion, and $c_{k}$ and $|\psi_{k}\rangle$ are the $k^{\textrm{th}}$ coefficient and $N$-electron SD, respectively. Generally called a CI wavefunction, the choice of the set of SDs results in approximate eigenstates for the Hamiltonian in $r$ spin-orbitals or qubits. For example, full CI (FCI) uses the expansion that includes all $N$-electron basis functions, and can represent the exact eigenstates of the Hamiltonian [65]. Such an approach scales combinatorially in $r$ and $N$, and polynomially-scaling approaches based on SCI are frequently used in classical simulations for computing many-electron ground states [15, 29]. When the wavefunction has several large CI coefficients, the state is commonly termed multireference; the purpose of the present work is quantum state preparation of multireference CI wavefunctions.
The algorithm has two main steps: superposition preparation, and Ansatz application, where we assume the quantum circuit is initially composed of $r$ qubits. The $r$-qubit Hilbert space contains many more Slater determanints than those in the $N$-electron space, but we must prepare the uniform superposition only over the selected determinants, which could be the FCI expansion or an unstructured subspace. For the Ansatz preparation, we apply a diagonal operator of CI coefficients to the uniform superposition to prepare the wavefunction. This diagonal operator is in general not unitary, so we use an ancillary qubit and construct a dilated diagonal unitary.
For preparing a superposition, the difference between a structured or unstructured subspace is important. We prepare the space of number-conserving SDs, known as the Dicke space [19], with a quantum circuit with CNOT-gate scaling of $\mathcal{O}(rN)$ and no ancilla qubits [47, 14]. We also prepare states which preserve both number and spin, as well as unstructured states, using a quantum walk algorithm that has CNOT-gate scaling of $\mathcal{O}(rD)$ [26]. The superposition over the entire $r$-qubit space requires no CNOT gates, and instead simply requires a Hadamard gate on each qubit. However, this space includes determinants which do not correspond to $N$-electron basis functions. We denote the subspace preparation as $U_{S}$,
$|S\rangle=U_{S}|0\rangle^{\otimes r},$ (3)
where $|S\rangle$ is the superposed subspace, and $|0\rangle^{\otimes r}$ indicates the $r$-fold tensor product of the ground state.
As an example, we consider the total uniform superposition over all $2^{r}$ states, meaning $U_{S}=H^{\otimes r}$. For coefficients $c_{k}$, we define a unitary matrix,
$U$ $=\begin{pmatrix}\Sigma_{+}&0\\ 0&\Sigma_{-}\end{pmatrix};\quad\Sigma_{\pm k}=c_{k}\pm i\sqrt{\frac{1-c_{k}}{\|c_{k}\|}}.$ (4)
We emphasize that $2\Sigma=\Sigma_{+}+\Sigma_{-}$}, which is a diagonal matrix with $c_{k}$ on the diagonal [59]. After applying $U_{S}$, we apply the Hadamard gate to the ancilla, followed by $U$, which acts over all $r+1$ qubits. Finally, we apply a Hadamard to the ancilla again. When the ancilla is measured in state $|0\rangle$ we prepare,
$\frac{1}{2}(\Sigma_{+}+\Sigma_{-})|S\rangle=\frac{1}{2}(2\Sigma)|S\rangle=|\Psi\rangle,$ (5)
and when the ancilla is in state $|1\rangle$, the negative combination of $\Sigma_{\pm}$ is prepared. Each of the wavefunctions conditioned on the ancilla outcome have unit norm, implying each outcome will occur exactly half of the time.
The dilated diagonal unitary $U$ can be implemented with Walsh operators as,
$U=\prod_{j~\textrm{odd}}e^{ia_{j}\hat{w}_{j}},$ (6)
where $a_{j}$’s are the Walsh coefficients, $\hat{w}_{j}$ are the $j^{\textrm{th}}$ order Walsh operators, and the product is over the entire system and ancilla space, $2^{r+1}$ [25]. Because $\Sigma_{+}$ and $\Sigma_{-}$ are complex conjugates, only half of the Walsh operators are needed. This operator can be implemented on a quantum circuit with CNOT and $R_{Z}$ gates using the binary expansion of $j$, most efficiently ordered by the Gray code [47, 72]. A schematic of the entire circuit is shown in Figure 1.
Figure 1: Circuit diagram for the preparation of $|\Psi\rangle$, where H is the Hadamard gate and $|a_{W}\rangle$ is the ancilla qubit for the Walsh Ansatz. The subspace selection step is performed either with the Dicke or quantum walk state preparation algorithm.
We express the Walsh coefficients $a_{j}$ directly in terms of the wavefunction coefficients $c_{k}$,
$a_{j}=2^{-r}~\sum_{k=1}^{2^{r}}c_{k}(-1)^{\langle k.j\rangle},$ (7)
where $\langle k.j\rangle$ is the binary inner product of the integer representations of the $k^{th}$ SD and Walsh order $j$. As noted above, $N$-electron wavefunctions do not span the complete $2^{r}$ Hilbert space, so the sum in Eq. 7 must be restricted to only those $D$ coefficients which are present in the chosen wavefunction expansion in Eq. 2. This also requires that the subspace superposition is prepared only over the $D$ chosen SDs.
Because the Walsh coefficients $a_{j}$ are the variational parameters of the quantum circuit, it is essential to reduce their number as well. An arbitrary selection of $D$ Walsh functions will not generally result in a full-rank Walsh-Fourier transform (WFT) for a $D$-dimensional wavefunction. To overcome this, we can oversample the Walsh basis by selecting $\mathcal{O}(D\textrm{log}D)$ functions randomly from the uniform distribution, which generates a full-rank WFT with probability scaling inversely with system size, namely $\mathcal{O}(1-D^{-1})$ [69]. The randomly selected Walsh functions $\hat{w}_{j}$ are then used in Eq. 6 and coefficients $a_{j}$ found with Eq. 7. This results in an overparameterized circuit, because there are $\mathcal{O}(D\textrm{log}D)$ circuit parameters and only $D$ wavefunction parameters, but this limits the classical preprocessing cost to matrix-vector multiplication, or $\mathcal{O}(D^{2}\textrm{log}D)$. In benchmark NISQ examples, we can also find $D$ orthogonal Walsh functions from an oversampled set using the QR decomposition, which scales as $\mathcal{O}(D^{3})$ and preserves the surjectivity of the Ansatz [71]. We use both approaches in the examples that follow.
III Results
We find the ground states of various molecules using VQE with either FCI or SCI, and benchmark the simulations with classical FCI and coupled-cluster single-doubles with perturbative triples (CCSD(T)) using PySCF [65, 13, 64]. We obtain the electronic Hamiltonian from OpenFermion, transformed to the qubit basis using the Jordan-Wigner transformation [73, 42], and use Qiskit to construct the quantum circuits for the statevector simulator, quantum device simulator, and IBMQ’s Torino quantum hardware, with details specified in the Appendix A [64, 30, 3]. To estimate the energy we use the Hadamard test to measure the expectation values of Pauli strings, requiring another ancilla qubit [4]. For H2, we converge the energy to chemical accuracy, $1.6\times 10^{-3}$ Hartree, and H6 we converge the energy to $10^{-5}$ Hartree, and use the COBYLA and BFGS optimizers implemented in SciPy respectively [71]. For the device and noisy simulator we sample with 1024 measurements. We find a full rank WFT by oversampling the Walsh basis and use the QR decomposition to find a $D$ by $D$ WFT.
In Figure 2 we show the the dissocation curves of the quantum FCI solutions for H2 and H6 in a minimal STO-6G basis [28].
Figure 2: Dissociation of (a) H2 using the subspace selected Walsh approach on IBMQ’s Torino processor (blue circles), Torino noisy simulator (green triangles) each with 1024 shots, and the FCI solution (line) and (b) linear H6 with a statevector simulator (teal squares), CCSD(T) (orange triangles), and FCI (black line). The red dashed line denotes 1.6 mHa error, which is chemical accuracy.
For H2 we use quantum walk to prepare the superposition, conserving spin and particle-number symmetries, which results in two SDs for the FCI wavefunction, namely $|1100\rangle$ and $|0011\rangle$ [65]. Figure 2 (a) shows the dissociation of H2 computed with FCI (black line), the Torino noisy simulator (blue circles), and the Torino quantum device (green triangles). The device and the simulator data are in excellent agreement with the FCI solution, almost always surpassing chemical accuracy. The dissociation curve of H6 is shown in Fig. 2 (b). We compare the results of the statevector simulator (teal squares), CCSD(T) (orange triangles), and FCI (black line). Here we include all spin preserving $N$-electron SDs, resulting in 400 parameters or SDs, and prepare the superposition with quantum walk. The ideal statevector simulator results highlight the accuracy of the present algorithm.
For large-scale applications and molecules in non-minimal basis sets, it is advantageous to avoid the classical QR decomposition. As described above we randomly oversample the Walsh basis with $D\textrm{log}D$ functions. This tradeoff between the surjectivity of the Ansatz and classical resource efficiency provides a practical approach for systems which are not possible with a QR decomposition. As an example we use H2O in a 6-31G basis set resulting in 44 circuit parameters for 16 SDs, shown in Figure 3 [20]. We show results from the statevector simulator (teal squares) and noiseless Aer sampler (purple diamonds) using 10 qubits and $2^{18}$ shots.
Figure 3: Dissociation of H2 in 6-31G basis with a statevector simulator (teal squares), noiseless Aer simulator (purple diamonds), using an oversampled Walsh basis. The red dashed line denotes 1.6 mHa error, which is chemical accuracy.
We also investigate whether all $D\textrm{log}D$ oversampled Walsh functions are necessary for a full rank transformation. Figure 4 shows the likelihood of finding a full rank transformation by random selection as a function of the fraction of $D\textrm{log}D$ for several molecules. To generate the data, we perform 50 random selections of Walsh functions from the uniform distribution and find the rank of the transform, and compute the success probability as the computed rank divided by the full rank. We find that for all molecules we can find a full rank transformation by random selection at about 50% of $D\textrm{log}D$. As the system size increases, a smaller fraction of $D\textrm{log}D$ Walsh functions is required to achieve a full rank transformation, which agrees with the previously reported $\mathcal{O}(1-D^{-1})$ results [69]. As expected, as system size increases, the required oversampling rate becomes smaller.
Figure 4: The median success probability of obtaining a full-rank WFT for several molecules, using fractions of $D\textrm{log}D$ randomly selected Walsh functions. Error bars represent the standard deviation of 50 trials.
We investigate the gate scaling of our quantum FCI approach by considering a series of molecules with $N$ electrons, $D$ SDs, and $r$ qubits. The subspace selection is achieved by either Dicke or quantum walk state preparation, whose CNOT gate scalings are $\mathcal{O}(rN)$ and $\mathcal{O}(rD)$, respectively. In Table 1 we show the number of CNOT gates for both the state preparations and the Walsh components with H6 and H8 chains, along with LiH, BeH2 and NH3 molecules.
Symmetry H6 (12 qubits) H8 (16 qubits) LiH (12 qubits) BeH2 (14 qubits) NH3 (16 qubits)
Spin & number 52 + 726 = 778 391 + 8756 = 9147 96 + 520 = 616 108 + 2328 = 2436 218 + 7530 = 7748
Number 382 + 1622 = 2004 723 + 22652 = 23375 265 + 1154 = 1419 476 + 6292 = 6768 840 + 18334 = 19174
No symmetry 0 + 4118 = 4118 0 + 65566 = 65566 0 + 4118 = 4118 0 + 16410 = 16410 0 + 65566 = 65566
Table 1: Number of CNOT gates for several molecules, where each row imposes different subspace symmetries. The first and the second numbers are the CNOT gate count for subspace selection and the Walsh Ansatz, respectively. Spin and particle-number conservation is enforced by the quantum walk algorithm; particle number conservation alone employs the Dicke state preparation algorithm. We show the number of system qubits, which is supplemented by two ancilla qubits.
In general, our resource requirements are similar to those of recent high-quality ADAPT-VQE circuits [6, 54]. Here, we find the exact FCI solution under three different cases: spin and particle-number conserving, particle-number conserving, and no symmetry.
The Walsh Ansatz has favorable scaling in both the two-qubit CNOT gates, and the number of variational parameters required for the one-qubit rotation gates with respect to the size of the parameter space. The number of CNOT gates for the Walsh Ansatz scales linearly with the number of determinants chosen, $\mathcal{O}(D)$. Furthermore, the Walsh Ansatz has the same number of variational parameters for $R_{Z}$ gates as the number of SDs in the selected expansion. This highlights that the Ansatz introduces no overparameterization of the chosen subspace, which helps in limiting spurious local minima in the optimization.
If we choose to represent the FCI wavefunction then the circuit will require exponentially scaling parameters and CNOT gates. Instead, we may select physically relevant CI coeffients using a variety of procedures. For example, starting from a classically computed CI expansion we can include excitations connected through a given order of perturbation theory or through a heuristic. The former procedure is standard in coupled-cluster and multireference perturbation theories [13, 7], and the latter is standard in classical SCI approaches [70, 60, 29].
We adapt the SCI approach to our quantum algorithm, using H2O in an STO-6G basis with a classically computed complete active space configuration interation (CASCI) wavefunction composed of 8 electrons in 10 spin orbitals [65]. From this wavefunction, we retain CI coefficients above a threshold $\epsilon$, then add SDs corresponding to single and double excitations from this truncated reference, which constitutes the variational wavefunction for the Walsh VQE. In this case we prepare the number- and spin-preserving subspace with the quantum walk algorithm. Table 2 shows the results for H2O at the equilibrium geometry, $r_{\textrm{OH}}=0.96$ Å and $\theta_{\textrm{OHO}}=104.3^{\circ}$, with two different $\epsilon$ thresholds and quantum FCI using a statevector simulator. We list the number of SDs, the number of CNOT gates, the resulting energy, and the fidelity compared to the exact solution. As expected, we see systematic improvement in the fidelity as the SCI space is increased.
H2O (14 qubits) SCI ($\epsilon=10^{-3}$) SCI ($\epsilon=10^{-7}$) FCI
# of SDs 227 267 441
# of CNOTs (SS) 579 378 60
# of CNOTs (W) 722 784 1352
Energy Error (Ha) 0.031 0.020 0.000
Fidelity 0.991 0.994 1.000
Table 2: Number of Slater determinants with two different thresholds for the SCI and quantum FCI, number of CNOT gates, energy error, and fidelity with FCI for H2O at its equilibrium geometry. ($S$) and (W) are quantum walk subspace selection and Walsh Ansatz, respectively. The quantum FCI result matches the classical result to machine precision. We show the number of system qubits, which needs to be supplemented by two ancilla qubits.
IV Conclusions and Discussion
Here, we present a variational quantum algorithm to estimate the ground-state energies of molecular Hamiltonians. This algorithm has two key steps, the first is preparing a state of physically-relevant bitstrings and the second is transforming the amplitudes of each SD with the truncated set of Walsh operators. After selecting the $D$ desired CI coefficients, we increase the parameter space to generate a full-rank transformation. A full-rank transformation can be generated by including $D\textrm{log}D$ randomly selected Walsh functions for the partial WFT [69]. The wavefunction Ansatz can be prepared with $\mathcal{O}(D\textrm{log}D)$ CNOT gates, and a $\mathcal{O}(D^{2}\textrm{log}^{2}D)$ classical operation. Furthermore, the Ansatz does not form a 2-design because it is diagonal and commutative, which mitigates the emergence of barren plateaus [41, 53, 52].
The subspace preparation step of the algorithm is very flexible. The only requirement is that a uniform superposition is created between the bitstrings of the subspace of interest, while all other amplitudes are zero. Through subspace selection, one can impose other symmetries such as parity or translational symmetry in condensed matter models [5, 8]. Effective application of symmetries remains an important tool in classical many-body methods, and our subspace selection protocol can act to symmetry-adapt wavefunctions in quantum algorithms [24, 35, 61, 46].
We discussed two ways to map CI coefficients to Walsh coefficients. We can preserve surjectivity through using a QR decomposition to find a full rank transformation, which has a maximum classical cost scaling $\mathcal{O}(D^{3})$ [58, 17, 11, 16]. The scaling of the QR decomposition is similar to diagonalization, so the classical cost of our algorithm using QR is similar to QSCI. QSCI diagonalizes a subspace CI Hamiltonian but is not iterative, while our algorithm, like all VQEs, is sensitive to the details of the iterative optimization [22, 55, 68]. For large-scale simulations, we can avoid the QR decomposition by relaxing the surjectivity requirement, and oversampling the Hilbert space by $\mathcal{O}(D\textrm{log}D)$, at a log-linear increase in the number of CNOT gates and parameters. This approach provides a scalable VQE strategy for both weakly- and strongly-correlated ground states.
SCI wavefunctions can also be constructed by alternative means, for example, by heat-bath or stochastic CI, or other heuristics [29, 77, 56]. Another possibility is preparing a high-quality guess wavefunction with the Walsh Ansatz, followed by application qubit or fermionic cluster operators [41, 67, 74, 6, 54], resulting in quantum formulations of multireference coupled cluster and perturbation theories [13, 7]. Taken together, our combined subspace preparation and Walsh function sampling approach provides a scalable and systematically improvable approach to quantum simulation of many-body ground states.
V Acknowledgements
KHM acknowledges start-up funding from the University of Minnesota. ARS acknowledges a National Science Foundation (NSF) REU summer fellowship (CHE-2349246). All authors acknowledge the computational resources provided by the Minnesota Supercomputing Institute (MSI) at the University of Minnesota. The authors acknowledge the use of IBM Quantum Credits for this work.
VI Code and Data Availability
The data from all results, along with the Python code to generate the data, are available at the Zenodo data repository [10].
Appendix A Appendix: Device detail
[…]